# GMSR:Gradient-Guided Mamba for Spectral Reconstruction from RGB Images

### v1

@@ III-D Spectral Gradient Attention
Fig.2(e) illustrates the spectral gradient attention. Define $F_{1}\in\mathbb{R}^{H\times W\times C}$ as the input feature map. Firstly, compute the gradient map along the spectral dimension using the first-order finite difference method, where the calculation is formulated as:
@@ III-E Loss Function
Method Params(M) FLOPS(G) NTIRE2020-CLEA CAVE Harvard
@@
MSCNN 26.6866 155.4068 0.0177 36.9092 0.9860 2.3036 0.0530 27.0281 0.8617 9.9603 0.0017 29.9853 0.9381 10.3051
HRNet 3.1705 624.5023 0.0152 38.7003 0.9924 1.7965 0.0541 28.1879 0.9425 7.8588 0.0016 30.7942 0.9594 7.6286
HSRnet 0.6826 160.5390 0.1619 39.1257 0.9869 2.0972 0.0568 27.0178 0.8910 9.5418 0.0015 31.0960 0.9633 7.4106
AGD-Net 0.3861 387.2278 0.0154 40.1399 0.9899 1.7361 0.0550 28.0085 0.9191 7.6613 0.0019 27.7370 0.8855 14.4955
MST++ 1.6196 88.8684 0.0162 38.0922 0.9910 1.7010 0.0740 24.6099 0.9218 13.098 0.0018 30.0349 0.9534 9.9057
RepCPSI 2.0778 511.6173 0.0165 41.3243 0.9919 1.6551 0.0469 29.8896 0.9566 6.1905 0.0016 30.9306 0.9546 10.0971
@@
Figure 4: Reconstruction results of different methods on the NTIRE2020 Clean dataset. The first row is ARAD_0463 (520 nm), the second row is ARAD_0457 (490 nm).
@@
Figure 7: Spectral response curves of randomly selected several spatial points from the reconstructed HSI of each SOTA method and the groundtruth HSI. (a) and (b) for the NTIRE2020 CLEAN. (c) and (d) for the CAVE. (e) and (f) for the Harvard.
@@ IV-A1 Datasets
To evaluate the performance of GMSR-Net, we conducted extensive experiments covering three widely used public SR datasets, including NTIRE2020clean [41], CAVE [42], and Harvard [43].
NTIRE2020 Clean: The NTIRE2020 Clean dataset is a sub-dataset of the NTIRE2020 Spectral Reconstruction Challenge, designed specifically for evaluating and comparing SR algorithms. The dataset has a spatial resolution of 512 × 482 and covers the wavelength range from 400 to 700 nm, with bands taken at 10 nm intervals, resulting in a total of 31 bands. The NTIRE2020 Clean dataset consists of 450 training images, 10 validation images, and 20 test images. Since the test set lacks HSIs corresponding to the RGB images, we use the validation set as an independent test set.
@@
Harvard: The Harvard dataset consists of 50 HSIs, each containing 31 bands. The spectral range covers 420-720nm, with a spatial resolution of 1392×1040. Similar to the CAVE dataset, we partition the dataset based on the alphabetical order of image names, with the first 43 images used for training and the remaining 7 images used for testing.
@@ IV-A3 Implementation Details
During the training process, we segmented the original RGB image into patches of size $128\times 128$. GMSR-Net used the Adam optimizer, where ${\beta_{1}}=0.9$, ${\beta_{1}}=0.99$, and $\epsilon=10-8$. For better training, we used a learning rate decay strategy, which initialized the learning rate to 0.0001, and set the power = 1.5. The implementation of GMSR-Net is based on the PyTorch framework and was trained on a single NVIDIA 4090Ti GPU for 100 epochs.
@@
Mamba Transformer CNN $G_{spa}$ $G_{spe}$ CBAM RMSE PSNR ASSIM SAM
@@
Transformer+GG 0.0150 40.8848 0.9908 1.7452
CNN+GG 0.0193 38.9331 0.9806 2.0876
Mamba+CBAM 0.0618 39.8417 0.9910 1.6303
Mamba+$G_{spa}$ 0.0154 40.4619 0.9911 1.6610
Mamba+$G_{spe}$ 0.0156 40.8263 0.9914 1.5110
@@
IV-B Comparison With SOTA Methods
To better demonstrate the effectiveness of the proposed GMSR-Net, we compared it with six state-of-the-art (SOTA) methods, namely MSCNN [44], HRNet [10], HSRnet [45], AGD-Net [46], MST++ [47], and RepCPSI [48].
Tab.I demonstrates the quantitative results of the different methods. It can be seen that the proposed GMSR achieves the best results in most of the metrics, which fully demonstrates the powerful generalization performance of the model. On the 2020NTIRE-CLEAN and Cave datasets, the SAM metrics of the GMSR method are reduced by 6.80% and 6.09%, respectively, which fully demonstrates the excellent performance of GMSR in spectral reconstruction. Additionally, thanks to the adoption of the spatial gradient attention module, GMSR achieved the best PSNR results on all three datasets, demonstrating excellent spatial reconstruction capabilities.
IV-C Model Efficiency Analyses
Tab.I shows the FLOPS and parameters of different methods at $512\times 482$ input size. The results show that the proposed GMSR-Net achieves efficient spectral reconstruction. Compared with the well-known HRNet, GMSR-Net drastically reduces the number of parameters and FLOPS by a factor of 70 and 40, respectively, while obtaining better reconstruction results. Compared to the latest RepCPSI, this reduction reaches 46 and 128 times, respectively. This demonstrates that GMSR-Net provides a powerful and promising lightweight backbone model for spectral reconstruction tasks, highlighting the potential of Mamba in driving model lightweight.
IV-D Ablation Studies
In order to analyze the role of each component of the GMSR-Net, we performed ablation studies on the NTIRE2020 Clean dataset. The ablation study is divided into two parts, an ablation study for each component of the network and another one for the number of Gradient Mamba blocks.
IV-D1 Effects of Different Ingredients
To validate the effectiveness of each component of our proposed GMSR-Net, we conducted the corresponding experiments and show the results in Tab.I. It is clear from the results that the reconstruction results show different degrees of degradation when we use either the Transformer or the CNN instead of Mamba block. In particular, when we remove the Mamba module in the network, the reconstruction results are significantly degraded. This further demonstrates the excellent characterization ability of Mamba in the reconstruction process, which provides critical support for our model performance. When we replaced $G_{spa}$ and $G_{spe}$ with the spatial and channel attention in CBAM respectively, we observed a significant decrease in all four metrics. Furthermore, removing $G_{spa}$ and $G_{spe}$ separately resulted in deteriorated reconstruction effects. Importantly, simultaneous removal of $G_{spa}$ and $G_{spe}$ led to a significant decrease in model performance. This result fully illustrates the importance of $G_{spa}$ and $G_{spe}$, which play a key role in maintaining the spatial spectral details and have a significant impact on improving the reconstruction results. Fig.8 illustrates the visual results of the above experiments, which are consistent with the quantitative results.
@@ IV-D2 Explore the Number of Gradient Mamba blocks
GMSR-Net consists of multiple GM blocks cascaded together. To investigate the effect of the number of GM blocks on the reconstruction results, we performed an ablation study. Tab.III and Fig.9 show the experimental quantitative and visual results. Defining n as the number of GM blocks, the experimental results show that as n increases, the reconstruction results become better and then worse. When the number of GM blocks is 3, the model achieves a good balance between accuracy and efficiency. Therefore, we finally chose to set the number of GM blocks to 3.
@@ V Conclusion
This paper introduces GMSR-Net, a novel lightweight spectral reconstruction network that leverages the Mamba architecture to model long-range dependencies. GMSR-Net incorporates spatial and spectral gradient attention mechanisms to enhance detail recovery during the reconstruction process. Our study underscores the potential of the Mamba model in addressing spectral reconstruction challenges and emphasizes the significance of gradient information in this process. The proposed GMSR-Net offers a powerful and promising lightweight solution for spectral reconstruction. In the future, we will further explore optimizing and extending this framework for broader applications in spectral reconstruction and related fields.

### v2

@@ III-D Spectral Gradient Attention
Fig. 2(e) illustrates the spectral gradient attention design. Define $F_{1}\in\mathbb{R}^{H\times W\times C}$ as the input feature map. Firstly, compute the gradient map along the spectral dimension using the first-order finite difference method, where the calculation is formulated as:
@@ III-E Loss Function
Method Params(M) FLOPs(G) NTIRE2020-Clean CAVE Harvard
@@
MSCNN [43] 26.6866 155.4068 0.0177 36.9092 0.9860 2.3036 0.0530 27.0281 0.8617 9.9603 0.0017 29.9853 0.9381 10.3051
HRNet [10] 3.1705 624.5023 0.0152 38.7003 0.9924 1.7965 0.0541 28.1879 0.9425 7.8588 0.0016 30.7942 0.9594 7.6286
HSRnet [44] 0.6826 160.5390 0.1619 39.1257 0.9869 2.0972 0.0568 27.0178 0.8910 9.5418 0.0015 31.0960 0.9633 7.4106
AGD-Net [45] 0.3861 387.2278 0.0154 40.1399 0.9899 1.7361 0.0550 28.0085 0.9191 7.6613 0.0019 27.7370 0.8855 14.4955
MST++ [46] 1.6196 88.8684 0.0162 38.0922 0.9910 1.7010 0.0740 24.6099 0.9218 13.098 0.0018 30.0349 0.9534 9.9057
RepCPSI [12] 2.0778 511.6173 0.0165 41.3243 0.9919 1.6551 0.0469 29.8896 0.9566 6.1905 0.0016 30.9306 0.9546 10.0971
@@
Figure 4: Reconstruction results comparison on the NTIRE2020-Clean. The first row is ARAD_0463 (520 nm) and the second row is ARAD_0457 (490 nm).
@@
Figure 7: Spectral response curves of randomly selected several spatial points from the reconstructed HSI of each SOTA method and the groundtruth HSI. (a) and (b) for the NTIRE2020-Clean. (c) and (d) for the CAVE. (e) and (f) for the Harvard.
@@ IV-A1 Datasets
To evaluate the performance of GMSR-Net, we conducted extensive experiments covering three widely used public SR datasets, including NTIRE2020-Clean [47], CAVE [48], and Harvard [49].
NTIRE2020-Clean: The NTIRE2020-Clean dataset is a sub-dataset of the NTIRE2020 Spectral Reconstruction Challenge, designed specifically for evaluating and comparing SR algorithms. The dataset has a spatial resolution of 512 × 482 and covers the wavelength range from 400 to 700 nm, with bands taken at 10 nm intervals, resulting in a total of 31 bands. The NTIRE2020-Clean dataset consists of 450 training images, 10 validation images, and 20 test images. Since the test set lacks HSIs corresponding to the RGB images, we use the validation set as an independent test set.
@@
Harvard: The Harvard dataset consists of 50 HSIs, each containing 31 bands. The spectral range covers 420-720nm, with a spatial resolution of 1392×1040. Similar to the CAVE dataset, we partition the dataset based on the alphabetical order of image names, with the first 46 images used for training and the remaining 4 images used for testing.
@@ IV-A3 Implementation Details
During the training process, we segmented the original RGB image into patches of size $128\times 128$. GMSR-Net used the Adam optimizer, where ${\beta_{1}}=0.9$, ${\beta_{1}}=0.99$, and $\epsilon=10^{-8}$. For better training, we used a learning rate decay strategy, which initialized the learning rate to 0.0001, and set the power = 1.5. The implementation of GMSR-Net is based on the PyTorch framework and was trained on a single NVIDIA 4090Ti GPU for 100 epochs.
IV-B Comparison With SOTA Methods
To better demonstrate the effectiveness of the proposed GMSR-Net, we compared it with six methods, namely MSCNN [43], HRNet [10], HSRnet [44], AGD-Net [45], MST++ [46], and RepCPSI [12].
Tab. I presents the quantitative results obtained by various methods. It is evident from the table that our proposed GMSR-Net consistently outperforms other methods across most of the evaluation metrics. This outcome underscores the robust generalization capabilities of our model. On the 2020NTIRE-CLEAN and Cave datasets, the SAM metrics of the GMSR-Net method are reduced by 6.80% and 6.09%, respectively, which fully demonstrates the excellent performance of GMSR-Net in spectral reconstruction. Additionally, thanks to the adoption of the spatial gradient attention module, GMSR-Net achieved the best PSNR results on all three datasets, demonstrating excellent spatial reconstruction capabilities.
IV-C Model Efficiency Analyses
Tab. I shows the FLOPs and parameters of different methods at $512\times 482$ input size. The results show that the proposed GMSR-Net achieves an excellent trade-off between speed and accuracy in SR task. Compared to the second-best performing method, RepCPSI, GMSR-Net reduces the number of parameters and FLOPs by an astounding 46.17×46.17$\times$46.17 × and $46.17\times 128.19$, respectively. In comparison with AGD-Net, which has the second fewest parameters, GMSR-Net reduces the parameter count by $46.17\times 8.58$. Similarly, it reduces the FLOPs by $46.17\times 5.99$ compared to MST++, the method with the second lowest FLOPS. This improvement is largely due to Mamba’s linear computational complexity when modeling long-range dependencies, allowing for higher efficiency in processing long sequences, especially under limited computational resources. Furthermore, GMSR-Net offers a powerful and promising lightweight backbone for SR, highlighting Mamba’s potential in driving model efficiency and lightweight design.
@@ IV-C Model Efficiency Analyses
VSS Transformer CNN $G_{spa}$ $G_{spe}$ CBAM RMSE PSNR ASSIM SAM
@@
Transformer+$G_{spa}$+$G_{spe}$ 0.0150 40.8848 0.9908 1.7452
CNN+$G_{spa}$+$G_{spe}$ 0.0193 38.9331 0.9806 2.0876
VSS+CBAM 0.0618 39.8417 0.9910 1.6303
VSS+$G_{spa}$ 0.0154 40.4619 0.9911 1.6610
VSS+$G_{spe}$ 0.0156 40.8263 0.9914 1.5110
@@
IV-D Ablation Studies
In order to analyze the role of each component of the GMSR-Net, we performed ablation studies on the NTIRE2020-Clean dataset. The ablation study is divided into two parts, an ablation study for each component of the network and another one for the number of GM blocks.
IV-D1 Effects of Different Ingredients
To validate the effectiveness of each component of our proposed GMSR-Net, we conducted the corresponding experiments and show the results in Tab. I. It is clear from the results that the reconstruction results show different degrees of degradation when we use either the Transformer or the CNN instead of VSS block. In particular, when we remove the VSS block in the network, the reconstruction results are significantly degraded. This further demonstrates the excellent characterization ability of Mamba in the reconstruction process, which provides critical support for our model performance. Define $G_{spa}$ and $G_{spe}$ as spatial gradient attention and spectral gradient attention block, respectively. When we replaced $G_{spa}$ and $G_{spe}$ with the spatial and channel attention in CBAM [50] respectively, we observed a significant decrease in all four metrics. Furthermore, removing $G_{spa}$ and $G_{spe}$ separately resulted in deteriorated reconstruction effects. Importantly, simultaneous removal of $G_{spa}$ and $G_{spe}$ led to a significant decrease in model performance. This result fully illustrates the importance of $G_{spa}$ and $G_{spe}$, which play a key role in maintaining the spatial-spectral details and have a significant impact on improving the reconstruction results. Fig. 8 illustrates the visual results of the above experiments. It can be seen that, compared to (b) and (i), the edges are significantly sharper, benefiting from the excellent edge sharpening capabilities of $G_{spa}$ and $G_{spe}$. Additionally, compared to (c) and (d), the overall error is noticeably smaller, demonstrating the outstanding global information recovery ability of the VSS block.
@@ IV-D2 Explore the Number of Gradient Mamba blocks
GMSR-Net consists of multiple GM blocks cascaded together. To investigate the effect of the number of GM blocks on the reconstruction results, we performed an ablation study. Tab. III and Fig. 9 show the experimental quantitative and visual results. By defining n as the number of GM blocks, our experimental findings reveal a trend wherein the quality of reconstruction initially improves with increasing n, but eventually deteriorates. Notably, when n is set to 3, the model strikes a favorable balance between accuracy and efficiency. This observation is further corroborated by Fig. 9, which visually depicts results consistent with our quantitative analyses. Consequently, we opt to fix the number of GM blocks at 3 for optimal performance.
@@ V Conclusion
This paper propose GMSR-Net, a novel lightweight spectral reconstruction model equipped with the long-range dependency-aware vision Mamba architecture. GMSR-Net subtly incorporates spatial and spectral gradient attention mechanisms to enhance detail recovery during the reconstruction process. Our study underscores the potential of the Mamba model in addressing spectral reconstruction challenges and emphasizes the significance of gradient information in this process. The proposed GMSR-Net offers a powerful and promising lightweight solution for spectral reconstruction. In the future, we will further extend this framework for broader applications in spectral reconstruction and related fields.